# Small maximal clusters are very unlikely in critical random graphs

### v1

@@ 1.4 Notation and structure of the paper
$\mathbb{P}\Big{(}\text{Bin}(n\lfloor BLT^{1/2}\rfloor,p)>np\lfloor BLT^{1/2}\rfloor+L(B^{2}/2-6)\Big{)}\leq\exp\big{(}-c\frac{L}{T^{1/2}}\big{)}\leq\exp\big{(}-cA^{3/2}\big{)}$ (24)
@@
$\mathbb{P}(t_{i}-t_{i-1}\geq T|\mathcal{F}_{i-1})\geq\frac{c_{0}}{T^{1/2}}$
for all large enough $n$, where we recall that $\mathcal{F}_{i-1}=\mathcal{F}_{t_{i-1}}$ is the $\sigma$-algebra consisting of all the information gathered by the exploration process until time $t_{i-1}$ (i.e. until the moment at which the procedure finishes revealing the $(i-1)$-th cluster). Then, setting $Y^{i}_{t}\coloneqq Y_{t_{i}+t}$ and $\eta^{i}_{t}\coloneqq\eta_{t_{i}+t}$ to simplify notation, we can write
$\mathbb{P}(t_{i}-t_{i-1}\geq T|\mathcal{F}_{i-1})=\mathbb{P}(Y^{i}_{t}>0\text{}\forall t<T|\mathcal{F}_{i-1})\geq\mathbb{P}\Big{(}1+\sum_{k=1}^{t}(\eta^{i}_{k}-1)>0\text{ }\forall t\leq T|\mathcal{F}_{i-1}\Big{)},$ (25)
where we recall that, setting $\mathcal{F}_{i-1}(k-1)\coloneqq\mathcal{F}_{t_{i-1}+k-1}$, we have
@@
Before proceeding with the actual proof, let us make a few observations and give some hints on how we intend to proceed.
Note that the $\eta^{i}_{k}$ are not independent, which makes our analysis non-trivial. The plan is to replace these random variables with a sequence of independent variables which are easier to analyse. To this end we first observe that, on $\mathcal{G}_{i-1}$, we have $t_{i-1}<BLT^{1/2}$; moreover, recalling that $L\sim An^{1/3},T\sim n^{2/3}/A$ and $A\ll n^{2/3}$ by assumption, we have $k\leq t<T\ll LT^{1/2}\ll n$. If we can show that $Y^{i}_{k-1}$ is at most $LT^{1/2}$ too, then we would obtain $t_{i-1}+k+Y^{i}_{k-1}\leq 3BLT^{1/2}$ and we could couple the $\text{Bin}_{k}(n-t_{i-1}-k+1-Y^{i}_{k-1},p)$ random variables with independent $\text{Bin}_{k}(n-3BLT^{1/2},p)$ random variables in such a way that
@@
thus removing the two sources of randomness from the first parameter of the (conditional) distribution in (26). At this stage, we would have a random walk $\sum_{k=1}^{t}[\text{Bin}_{k}(n-3BLT^{1/2},p)-1]$ with independent and identically distributed increments, having a small negative drift; the drift is so small that it does not impact much the behaviour of the process, which can be regarded as a random walk with iid, mean zero increments having finite second moment. Putting all pieces together, we would bound from below the expression in (25) by the probability that a mean-zero $\mathbb{Z}$-valued random walk with the latter properties stays above zero for $T$ steps, which is known to be $\asymp T^{-1/2}$ (as desired).
We now make rigorous the above simple strategy, and start by replacing the $\eta^{i}_{k}$ with iid random variables having the $\text{Bin}(n-3L\lceil BT^{1/2}\rceil,p)$ law.
@@
$\mathbb{P}\Big{(}1+\sum_{k=1}^{t}(\eta^{i}_{k}-1)>0\text{ }\forall t\leq T|\mathcal{F}_{i-1}\Big{)}\geq\mathbb{P}\Big{(}1+\sum_{k=1}^{t}(\delta_{k}-1)>0\text{ }\forall t\leq T\Big{)}-\exp(-cLT^{1/2})$
@@
Recall that $u_{k}$ is the vertex that is explored at step $t$ in the exploration of $\mathbb{G}(n,p)$. Denoting by $X^{v}_{k}$ the indicator that vertex $u_{k}$ is a neighbour of $v\in[n]$, we note that the random variable $\eta^{0}_{k}$ can be expressed as
$\eta^{i}_{k}=\sum_{v\in[n]\setminus(\mathcal{A}^{i}_{k-1}\cup\mathcal{E}^{i}_{k-1})}X^{v}_{k},$
with $\mathcal{E}^{i}_{j}\coloneqq\mathcal{E}_{t_{i-1}+j}$ being the set of explored vertices at time $t_{i-1}+j$. Let us denote by $\mathcal{R}^{i}_{j}\coloneqq\mathcal{A}^{i}_{j}\cup\mathcal{E}^{i}_{j}$ the set of vertices which are either active or explored at time $j$ (we use $\mathcal{R}$ for ‘revealed’). For convenience, we also give a name to the set of unseen vertices which become active at step $j$ of the exploration of $\mathcal{C}_{i-1}$ (i.e. at time $t_{i-1}+j$); denote such a set by $\mathcal{A}^{i,*}_{j}$, so that then $\mathcal{R}^{i}_{k-1}=\bigcup_{j=0}^{k-1}\mathcal{A}^{i,*}_{j}$ (with $\mathcal{A}^{i,*}_{0}=\{u_{t_{i-1}+1}\}$, the vertex from which we start the exploration of $\mathcal{C}_{i}$).
For each $1\leq k\leq T$, if $|\mathcal{R}^{i}_{k-1}|<3L\lceil BT^{1/2}\rceil$ (which occurs if $Y^{i}_{k-1}<BLT^{1/2}$, because $|\mathcal{R}^{i}_{k-1}|=Y^{i}_{k-1}+t_{i-1}+k-1<Y^{i}_{k-1}+2BLT^{1/2}$ on the event $\mathcal{G}_{i-1}$) we let $\mathcal{B}^{i,*}_{k-1}$ be any subset of the vertices $[n]$ such that
@@
$\delta_{k}\coloneqq\sum_{v\in[n]\setminus(\mathcal{B}^{i,*}_{k-1}\cup\mathcal{R}^{i}_{k-2})}X^{v}_{k}+\sum_{i=1}^{h_{k}}I^{k}_{j}.$
Note that $|[n]\setminus(\mathcal{B}^{i,*}_{k-1}\cup\mathcal{R}^{i}_{k-2})|=n-3L\lceil BT^{1/2}\rceil$ by construction and hence
@@
Moreover, the random variables $\delta_{k}$ are independent because $\{X^{k}_{v}:v\in\mathcal{(}R^{i}_{k-2})^{c}\}$ are independent and independent of $\{X^{j}_{v}:v\in\mathcal{(}R^{i}_{j-2})^{c}\}$ for any $j\neq k$; they are also independent of the indicators associated to the edges revealed up to time $t_{i-1}$, whence independent of $\mathcal{F}_{i-1}$.
We also observe that if $|\mathcal{R}^{i}_{k-1}|<3L\lceil BT^{1/2}\rceil$, then $|\mathcal{B}^{i,*}_{k-1}\cup\mathcal{R}^{i}_{k-2}|=3L\lceil BT^{1/2}\rceil$ and so $h_{k}=0$. Since we also have $\mathcal{A}^{i,*}_{k-1}\subset\mathcal{B}^{i,*}_{k-1}$ by construction, we see that if $|\mathcal{R}^{i}_{k-1}|<3L\lceil BT^{1/2}\rceil$ then $\eta^{i}_{k}\geq\delta_{k}$. Thus, recalling that $t_{i}$ denotes the first time at which the set of active vertices becomes empty for the first time after $t_{i-1}$, we see that the probability which appears in the statement of the lemma is at least
$\mathbb{P}\Big{(}1+\sum_{k=1}^{t}(\delta_{k}-1)>0\text{ }\forall t\leq T,Y^{i}_{k}<BLT^{1/2}\text{ }\forall k<(t_{i}-t_{i-1})\wedge T|\mathcal{F}_{i-1}\Big{)},$ (27)
where the last inequality follows from the fact that, on the event $\{Y^{i}_{k}<BLT^{1/2}\text{ }\forall k<T\}$, we have $|\mathcal{R}^{i}_{k-1}|<3L\lceil BT^{1/2}\rceil$ for each $k$ as discussed earlier, and so $\eta^{i}_{k}\geq\delta_{k}$ for $k<T$ (and $(t_{i}-t_{i-1})\wedge T=T$ on the event $\{1+\sum_{k=1}^{t}(\eta^{i}_{k}-1)>0\text{ }\forall t\leq T-1\}$). To conclude, we observe that the probability on the right-hand side of (27) is at least
$\mathbb{P}\Big{(}1+\sum_{k=1}^{t}(\delta_{k}-1)>0\text{ }\forall t\leq T\Big{)}-\mathbb{P}\Big{(}\exists k<(t_{i}-t_{i-1})\wedge T:Y^{i}_{k}\geq BLT^{1/2}|\mathcal{F}_{i-1}\Big{)},$
where we have used the above mentioned independence for the last inequality. There remains to bound from above the second probability on the right-hand side of the last expression, which is easily done. Indeed, using a union bound together with the fact that $Y^{i}_{k}=1+\sum_{j=1}^{k}(\eta^{i}_{j}-1)$ for $k\leq t_{i}-t_{j-1}$, we can write
$\mathbb{P}\Big{(}\exists k<(t_{i}-t_{i-1})\wedge T:Y^{i}_{k}\geq BLT^{1/2}|\mathcal{F}_{i-1}\Big{)}\leq\sum_{k=1}^{T-1}\mathbb{P}\Big{(}1+\sum_{j=1}^{k}(\eta^{i}_{j}-1)\geq BLT^{1/2}|\mathcal{F}_{i-1}\Big{)}.$
By coupling the $\eta^{i}_{k}$ with independent $\text{Bin}(n,p)$ random variables $\xi_{k}$ in such a way that $\eta^{i}_{k}\leq\xi_{k}$ and using Lemma 1.1, we see that the probability on the right-hand side of the last expression is at most
$\sum_{k=1}^{T-1}\exp\big{(}-c\frac{B^{2}i^{2}T}{knp+BiT^{1/2}}\big{)}\leq\exp\big{(}-cLT^{1/2}[1-\log(T)/(cLT^{1/2})]\big{)}\leq\exp\big{(}-cLT^{1/2}),$
where we have used that $\log(T)\leq T\ll LT^{1/2}$. ∎
@@
where $\hat{M}_{t}=1+\sum_{i=1}^{t}(\hat{W}_{i}-1)$, and $(\hat{W}_{i})_{i=1}^{t_{n}}$ is a sequence of independent Poisson random variables with mean one.
@@
Note that clearly $na_{n}\in\mathbb{Z}$ and, since $LT^{1/2}/n\sim A^{1/2}/n^{1/3}\ll 1$ (as we assumed that $A\ll n^{2/3}$), we also have $a_{n}>-1$ for all sufficiently large $n$. Moreover, it is clear that $b_{n}\in(-1,n-1)$ provided $n$ is large enough.
Now, setting $\mu_{n}\coloneqq(1+a_{n})(1+b_{n})$ as in the statement of Lemma 2.4, it is not difficult to see that $\mu_{n}$ is negative provided $A\geq A_{0}$ for a large enough $A_{0}=A_{0}(\lambda)$ which depends on $\lambda$ and moreover
$|1-\mu_{n}|=O\big{(}\frac{LT^{1/2}}{n}\big{)}\asymp\frac{A^{1/2}}{n^{1/3}}\ll 1.$
@@
for some positive constant $c=c(B)>0$ which depends solely on $B$. But then we can bound, using Lemma 2.4 (note that for us $t_{n}=T$ and $h_{n}=\lceil T^{1/2}\rceil$)
$\mathbb{P}\Big{(}\sum_{k=1}^{t}(\delta_{k}-1)$ $>0\text{ }\forall t\leq T\Big{)}$
$\geq\mathbb{P}\Big{(}\sum_{k=1}^{t}(\delta_{k}-1)>0\text{ }\forall t\leq T,\sum_{k=1}^{T}(\delta_{k}-1)\in[\lceil T^{1/2}\rceil,2\lceil T^{1/2}\rceil]\Big{)}$
$\geq c(B)\mathbb{P}\big{(}M_{t}>0\text{ }\forall t\leq T,M_{T}\in[\lceil T^{1/2}\rceil,2\lceil T^{1/2}\rceil]\big{)}-\frac{T}{2n}$ (29)
@@
$\mathbb{P}\big{(}M_{t}>0\text{ }\forall t\leq T,M_{T}\in[\lceil T^{1/2}\rceil,2\lceil T^{1/2}\rceil]\big{)}\geq c\sum_{j=\lceil T^{1/2}\rceil}^{2\lceil T^{1/2}\rceil}\frac{j}{T^{3/2}}\geq cT^{-1/2}.$ (31)
Finally, observe that the term $T/2n$ on the right-hand side of (2.2.1) is $O((An^{1/2})^{-1})$, which is much smaller than $A^{1/2}/n^{1/3}\asymp T^{-1/2}$.
Therefore we conclude that there exists a constant $c_{0}=c_{0}(B)>0$ which depends solely on $B$ such that, for all large enough $A,n$,
$\mathbb{P}\Big{(}\sum_{k=1}^{t}(\delta_{k}-1)>0\text{ }\forall t\leq T\Big{)}\geq c_{0}T^{-1/2},$
@@
We start by describing a method for exploring the components of the graph $\mathbb{G}(n,d,p)$, which we recall is the random graph obtained by performing bond percolation with parameter $p$ on a realisation $\mathbb{G}(n,d)$ of a $d$-regular graph sampled uniformly at random from the set of all $d$-regular (simple) graphs on $[n]$. The exploration process we use here is taken verbatim from [13] (see also [20] and [27]).
@@
for $i\geq 1$ (prior to the end of the procedure). The next result, which corresponds to Lemma $10$ in [27], gives us a way to obtain an upper bound for the probability that $\mathcal{C}_{\max}(\mathbb{G}^{\prime}(n,d,p))$ contains less than $n^{2/3}/A$ vertices in terms of the (random) lengths of the positive excursions of the process $Y_{t}$. We refer the reader to [10] for an intuitive explanation.
@@
As we did when analysing the $\mathbb{G}(n,d,p)$ model, we define the (good) event
$\mathcal{G}_{j}\coloneqq\{t_{j}\leq BLT^{1/2}\},$
where $L\leq j\leq 2L$ and $B$ is some large positive constant (independent of $A,n,\lambda$ but possibly dependent on $d$). The events $\mathcal{G}_{j}$ are needed to bound from above the number of explored half-edges while revealing the cluster of a vertex in $\mathbb{G}^{\prime}(n,d,p)$; such an upper bound on the number of explored stubs is needed in order to derive a meaningful lower bound for the probability that the number of active half-edges stays above zero for the required number of steps.
@@
$\mathbb{P}(t_{i}-t_{i-1}<T\text{ }\forall L+1\leq i\leq 2L)\leq\mathbb{P}(\{t_{i}-t_{i-1}<T\text{ }\forall L+1\leq i\leq 2L\}\cap\mathcal{G}_{2L})+\mathbb{P}(\mathcal{G}^{c}_{2L}).$
The next lemma controls the probability of the (bad) event $\mathcal{G}^{c}_{2L}$.
Lemma 2.6.
@@
Therefore in what follows we can focus on the probability
($\mathbb{P}(\{t_{i}-t_{i-1}<T\text{ }\forall L+1\leq i\leq 2L\}\cap\mathcal{G}_{2L})$ (35)
Let us define $\mathcal{F}_{j}\coloneqq\mathcal{F}_{t_{j}}$ for $j\in[N]$, the $\sigma$-algebra consisting of all the information collected by the exploration process by time $t_{j}$, i.e. by the time at which we start the exploration of the $j$-th component. Wit the next result we show that, on the event $\mathcal{G}_{i-1}$, we have $t_{i}-t_{i-1}\geq T$ with probability $\gtrsim T^{-1/2}$.
Lemma 2.7.
@@
$\mathbb{P}(t_{i}-t_{i-1}\geq T|\mathcal{F}_{i-1})\geq\frac{c_{0}}{T^{1/2}}$
@@
($\mathbb{P}(\{t_{i}-t_{i-1}<T\text{ }\forall L+1\leq i\leq 2L\}\cap\mathcal{G}_{2L})\\ \leq\mathbb{E}\Big{[}\mathbbm{1}_{\{t_{i}-t_{i-1}<T\text{ }\forall L+1\leq i\leq 2L-1\}\cap\mathcal{G}_{2L-1}}\big{(}1-\mathbb{P}(t_{2L}-t_{2L-1}\geq T|\mathcal{F}_{2L-1})\big{)}\Big{]}.$ (36)
Thanks to Lemma 2.7 we know that the expression on the right-hand side of (36) is at most
@@
Iterating and using the classical inequality $1+x\leq e^{x}$ (which is valid for all $x\in\mathbb{R}$), we conclude that

### v3

@@ 1.4 Some useful tools
$\mathbb{P}\Big{(}\text{Bin}(n\lfloor BLT^{1/2}\rfloor,p)>np\lfloor BLT^{1/2}\rfloor+L(B^{2}/12-4)\Big{)}\leq\exp\big{(}-c\frac{L}{T^{1/2}}\big{)}\leq\exp\big{(}-cA^{3/2}\big{)}$ (24)
@@
$\mathbb{P}(t_{i}-t_{i-1}\geq T|\mathcal{F}_{t_{i-1}})\geq\frac{c_{0}}{T^{1/2}}$
for all large enough $n$, where we recall that $\mathcal{F}_{t_{i-1}}$ is the $\sigma$-algebra consisting of all the information gathered by the exploration process until time $t_{i-1}$ (i.e. until the moment at which the procedure finishes revealing the $(i-1)$-th cluster). Setting $Y^{i}_{t}\coloneqq Y_{t_{i-1}+t}$ and $\eta^{i}_{t}\coloneqq\eta_{t_{i-1}+t}$ to simplify notation, we can write
$\mathbb{P}(t_{i}-t_{i-1}\geq T|\mathcal{F}_{t_{i-1}})=\mathbb{P}(Y^{i}_{t}>0\text{ }\forall t<T|\mathcal{F}_{t_{i-1}})\geq\mathbb{P}\Big{(}1+\sum_{k=1}^{t}(\eta^{i}_{k}-1)>0\text{ }\forall t\leq T|\mathcal{F}_{t_{i-1}}\Big{)},$ (25)
where we recall that, setting $\mathcal{F}_{i-1}(k-1)\coloneqq\mathcal{F}_{t_{i-1}+k-1}$,
@@
Before proceeding with the actual proof, let us make a few observations and give some hints on how we intend to proceed. (A similar heuristic was given in Section 2.1).
Note that the $\eta^{i}_{k}$ are not independent, which makes our analysis non-trivial. The plan is to replace these random variables with a sequence of independent variables which are easier to analyze. To this end we first observe that, on the event $\mathcal{G}_{i-1}$, we have $t_{i-1}<BLT^{1/2}$; moreover, recalling that $L\sim An^{1/3}/2,T\sim n^{2/3}/A$ and $A\ll n^{2/3}$ by assumption, we have $k\leq t<T\ll LT^{1/2}(\ll n)$. If we can show that $Y^{i}_{k-1}$ is at most $LT^{1/2}$ too, then we would obtain $t_{i-1}+k+Y^{i}_{k-1}\leq 3BLT^{1/2}$ and we could couple the $\text{Bin}_{k}(n-t_{i-1}-k+1-Y^{i}_{k-1},p)$ random variables with independent $\text{Bin}_{k}(n-3BLT^{1/2},p)$ random variables in such a way that
@@
thus removing the two sources of randomness from the first parameter of the (conditional) distribution in (26). At this stage, we would have a random walk with independent and identically distributed increments $\text{Bin}_{k}(n-3BLT^{1/2},p)-1$ having a small negative drift; the drift is so small that it does not impact much the behavior of the process, which can be regarded as a random walk with iid, mean zero increments having finite second moment. Putting all pieces together, we would bound from below the expression in (25) by the probability that a mean-zero $\mathbb{Z}$-valued random walk with finite variance stays above zero for $T$ steps, which is known to be $\Theta(T^{-1/2})$ (as desired).
We now make rigorous the above simple argument, and start by replacing the $\eta^{i}_{k}$ with iid random variables having the $\text{Bin}(n-3L\lceil BT^{1/2}\rceil,p)$ law.
@@
$\mathbb{P}\Big{(}1+\sum_{k=1}^{t}(\eta^{i}_{k}-1)>0\text{ }\forall t\leq T|\mathcal{F}_{t_{i-1}}\Big{)}\geq\mathbb{P}\Big{(}1+\sum_{k=1}^{t}(\delta_{k}-1)>0\text{ }\forall t\leq T\Big{)}-\exp(-cLT^{1/2})$
@@
Recall that $u_{k}$ is the vertex that is explored at step $k$ in the exploration of $\mathbb{G}(n,p)$. Denoting by $X^{i}_{k}(v)$ the indicator that vertex $u_{t_{i-1}+k}$ is a neighbor of $v\in[n]$, we note that the random variable $\eta^{i}_{k}$ can be expressed as
$\eta^{i}_{k}=\sum_{v\in[n]\setminus(\mathcal{A}^{i}_{k-1}\cup\mathcal{E}^{i}_{k-1})}X^{i}_{k}(v),$
with $\mathcal{E}^{i}_{j}\coloneqq\mathcal{E}_{t_{i-1}+j}$ and $\mathcal{A}^{i}_{j}\coloneqq\mathcal{A}_{t_{i-1}+j}$ being the sets of explored and active vertices at time $t_{i-1}+j$, respectively. Let us denote by $\mathcal{R}^{i}_{j}\coloneqq\mathcal{A}^{i}_{j}\cup\mathcal{E}^{i}_{j}$ the set of vertices which are either active or explored at time $t_{i-1}+j$ (we use the letter $\mathcal{R}$ for ‘revealed’). We also give a name to the set of unseen vertices which become active at step $j$ of the exploration of $\mathcal{C}_{i}$ (i.e. at time $t_{i-1}+j$); denote such a set by $\mathcal{A}^{i,*}_{j}$, so that then $\mathcal{R}^{i}_{k-1}=\bigcup_{j=0}^{k-1}\mathcal{A}^{i,*}_{j}$ (with $\mathcal{A}^{i,*}_{0}=\{u_{t_{i-1}+1}\}$, the vertex from which we start the exploration of $\mathcal{C}_{i}$).
For each $1\leq k\leq T$, if $|\mathcal{R}^{i}_{k-1}|<3L\lceil BT^{1/2}\rceil$ (which occurs when $Y^{i}_{k-1}<BLT^{1/2}$, because $|\mathcal{R}^{i}_{k-1}|=Y^{i}_{k-1}+t_{i-1}+k-1<Y^{i}_{k-1}+2BLT^{1/2}$ on the event $\mathcal{G}_{i-1}$) we let $\mathcal{B}^{i,*}_{k-1}$ be any subset of $[n]$ such that
@@
$\delta_{k}\coloneqq\sum_{v\in[n]\setminus(\mathcal{B}^{i,*}_{k-1}\cup\mathcal{R}^{i}_{k-2})}X^{i}_{k}(v)+\sum_{i=1}^{h_{k}}I^{k}_{j}.$
Note that
@@
Moreover, the random variables $\delta_{k}$ are independent because the $\{X^{i}_{k}(v):v\in(\mathcal{R}^{i}_{k-2})^{c}\}$ are independent and independent of $\{X^{i}_{j}(v):v\in(\mathcal{R}^{i}_{j-2})^{c}\}$ for any $j\neq k$; they are also independent of $\mathcal{F}_{t_{i-1}}$.
We also observe that, if $|\mathcal{R}^{i}_{k-1}|<3L\lceil BT^{1/2}\rceil$, then $|\mathcal{B}^{i,*}_{k-1}\cup\mathcal{R}^{i}_{k-2}|=3L\lceil BT^{1/2}\rceil$ and so $h_{k}=0$. Since we also have $\mathcal{A}^{i,*}_{k-1}\subset\mathcal{B}^{i,*}_{k-1}$ by construction, we see that if $|\mathcal{R}^{i}_{k-1}|<3L\lceil BT^{1/2}\rceil$ then $\eta^{i}_{k}\geq\delta_{k}$. Thus, recalling that $t_{i}$ denotes the first time at which the set of active vertices becomes empty after time $t_{i-1}$, we see that the probability which appears in the statement of the lemma is at least
$\mathbb{P}\Big{(}1+\sum_{k=1}^{t}(\delta_{k}-1)>0\text{ }\forall t\leq T,Y^{i}_{k}<BLT^{1/2}\text{ }\forall k<(t_{i}-t_{i-1})\wedge T|\mathcal{F}_{t_{i-1}}\Big{)}.$ (27)
Indeed, on the event $\{Y^{i}_{k}<BLT^{1/2}\text{ }\forall k<T\}$, we have $|\mathcal{R}^{i}_{k-1}|<3L\lceil BT^{1/2}\rceil$ for each $k$ as discussed earlier, and so $\eta^{i}_{k}\geq\delta_{k}$ for $k<T$ (and $(t_{i}-t_{i-1})\wedge T=T$ on the event $\{1+\sum_{k=1}^{t}(\eta^{i}_{k}-1)>0\text{ }\forall t\leq T-1\}$). To conclude, we observe that the probability on the right-hand side of (27) is at least
$\mathbb{P}\Big{(}1+\sum_{k=1}^{t}(\delta_{k}-1)>0\text{ }\forall t\leq T\Big{)}-\mathbb{P}\Big{(}\exists k<(t_{i}-t_{i-1})\wedge T:Y^{i}_{k}\geq BLT^{1/2}|\mathcal{F}_{t_{i-1}}\Big{)},$
where we have used the above mentioned independence of the $\delta_{k}$ and $\mathcal{F}_{t_{i-1}}$. There remains to bound from above the second probability on the right-hand side of the last expression, which is easily done. Indeed, using a union bound together with the fact that $Y^{i}_{k}=1+\sum_{j=1}^{k}(\eta^{i}_{j}-1)$ for $k\leq t_{i}-t_{j-1}$, we can write
$\mathbb{P}\Big{(}\exists k<(t_{i}-t_{i-1})\wedge T:Y^{i}_{k}\geq BLT^{1/2}|\mathcal{F}_{t_{i-1}}\Big{)}\leq\sum_{k=1}^{T-1}\mathbb{P}\Big{(}1+\sum_{j=1}^{k}(\eta^{i}_{j}-1)\geq BLT^{1/2}|\mathcal{F}_{t_{i-1}}\Big{)}.$
By coupling the $\eta^{i}_{j}$ with independent $\text{Bin}(n,p)$ random variables $\xi_{j}$ in such a way that $\eta^{i}_{j}\leq\xi_{j}$ and using Lemma 1.1 we see that, if $A$ is large enough, then the probability on the right-hand side of the last expression is at most
$\sum_{k=1}^{T-1}\exp\big{(}-c\frac{B^{2}L^{2}T}{knp+BLT^{1/2}}\big{)}\leq\exp\big{(}-cLT^{1/2}[1-\log(T)/(cLT^{1/2})]\big{)}\leq\exp\big{(}-cLT^{1/2})$
for some $c=c(B)>0$, where we have used that $\log(T)\leq T\ll LT^{1/2}$. ∎
@@
where $\hat{M}_{t}=1+\sum_{i=1}^{t}(\hat{W}_{i}-1)$, and $(\hat{W}_{i})_{i=1}^{t_{n}}$ is a sequence of independent $\text{Poi}(1)$ random variables.
@@
Note that clearly $na_{n}\in\mathbb{Z}$ and, since $LT^{1/2}/n\sim A^{1/2}/2n^{1/3}\ll 1$ (as we assumed that $A\ll n^{2/3}$), we also have $a_{n}>-1$ for all sufficiently large $n$. Moreover, it is clear that $b_{n}\in(-1,n-1)$ provided $n$ is large enough.
Now, setting $\mu_{n}\coloneqq(1+a_{n})(1+b_{n})$ as in the statement of Lemma 2.2, it is not difficult to see that $\mu_{n}$ is negative provided $A\geq A_{0}$ for a large enough $A_{0}=A_{0}(\lambda)>0$ which depends on $\lambda$ and moreover
$|1-\mu_{n}|=O\big{(}\frac{LT^{1/2}}{n}\big{)}=O\big{(}\frac{A^{1/2}}{n^{1/3}}\big{)}\ll 1.$
@@
for some positive constant $c=c(B)>0$ which depends solely on $B$. But then we can bound, using Lemma 2.2 (with $t_{n}=T$ and $h_{n}=\lceil T^{1/2}\rceil$)
$\mathbb{P}\Big{(}1+\sum_{k=1}^{t}(\delta_{k}-1)>0\text{ }\forall t\leq T\Big{)}$ $\geq\mathbb{P}\Big{(}1+\sum_{k=1}^{t}(\delta_{k}-1)>0\text{ }\forall t\leq T,1+\sum_{k=1}^{T}(\delta_{k}-1)\in[\lceil T^{1/2}\rceil,2\lceil T^{1/2}\rceil]\Big{)}$
$\geq c(B)\mathbb{P}\big{(}M_{t}>0\text{ }\forall t\leq T,M_{T}\in[\lceil T^{1/2}\rceil,2\lceil T^{1/2}\rceil]\big{)}-O\big{(}frac{T}{2n}\big{)}$ (29)
@@
$\mathbb{P}\big{(}M_{t}>0\text{ }\forall t\leq T,M_{T}\in[\lceil T^{1/2}\rceil,2\lceil T^{1/2}\rceil]\big{)}\geq c\sum_{j=\lceil T^{1/2}\rceil-1}^{2\lceil T^{1/2}\rceil-1}\frac{j}{T^{3/2}}\geq cT^{-1/2}.$ (31)
Finally, observe that the term $T/n$ on the right-hand side of (2.2.1) is $O((An^{1/3})^{-1})$, which is much smaller than $A^{1/2}/n^{1/3}=\Theta(T^{-1/2})$. Therefore we conclude that there exists a constant $c_{0}=c_{0}(B)>0$ which depends solely on $B$ such that, for all large enough $A,n$,
$\mathbb{P}\Big{(}1+\sum_{k=1}^{t}(\delta_{k}-1)>0\text{ }\forall t\leq T\Big{)}\geq c_{0}T^{-1/2},$
@@
We start by recalling a method for exploring the components of the graph $\mathbb{G}(n,d,p)$, which we recall is the random graph obtained by performing bond percolation with parameter $p$ on a realization $\mathbb{G}(n,d)$ of a $d$-regular graph sampled uniformly at random from the set of all $d$-regular (simple) graphs on $[n]$. The exploration process we use here is taken verbatim from [13] (see also [20] and [27]).
@@
for $i\geq 1$ (prior to the end of the procedure). The next result, which corresponds to Lemma $10$ in [27], gives us a way to obtain an upper bound for the probability that $\mathcal{C}_{\max}(\mathbb{G}^{\prime}(n,d,p))$ contains less than $n^{2/3}/A$ vertices in terms of the (random) lengths of the positive excursions of the process $Y_{t}$. (We refer the reader to [10] for an intuitive explanation.)
@@
As we did when analyzing the $\mathbb{G}(n,d,p)$ model, we define the (good) event
$\mathcal{G}_{j}=\mathcal{G}_{j}(B)\coloneqq\{t_{j}\leq BLT^{1/2}\},$
where $L\leq j\leq 2L-1$ and $B$ is some large positive constant (independent of $A,n,\lambda$ but possibly dependent on $d$). The events $\mathcal{G}_{j}$ are needed to bound from above the number of explored half-edges while revealing the cluster of a vertex in $\mathbb{G}^{\prime}(n,d,p)$; such an upper bound on the number of explored stubs is needed in order to derive a meaningful lower bound for the probability that the number of active half-edges stays above zero for the required number of steps.
@@
$\mathbb{P}(t_{i}-t_{i-1}<T\text{ }\forall L+1\leq i\leq 2L)\leq\mathbb{P}(\{t_{i}-t_{i-1}<T\text{ }\forall L+1\leq i\leq 2L\}\cap\mathcal{G}_{2L-1})+\mathbb{P}(\mathcal{G}^{c}_{2L-1}).$
The next proposition controls the probability of the (bad) event $\mathcal{G}^{c}_{2L-1}$ and shows that it is at most $\exp(-\Omega(A^{3/2}))$.
Proposition 2.3.
@@
Therefore, in what follows we can focus on the probability
($\mathbb{P}(\{t_{i}-t_{i-1}<T\text{ }\forall L+1\leq i\leq 2L\}\cap\mathcal{G}_{2L-1})$ (35)
Let us write $\mathcal{F}_{j}$ for the $\sigma$-algebra consisting of all the information collected by the exploration process by time $j$; then $\mathcal{F}_{t_{j}}$ is the information gathered by the time at which we finish the exploration of the $j$-th component. Wit the next result we show that, on the event $\mathcal{G}_{i-1}$, we have $t_{i}-t_{i-1}\geq T$ with probability at least of order $T^{-1/2}$.
Proposition 2.4.
@@
$\mathbb{P}(t_{i}-t_{i-1}\geq T|\mathcal{F}_{t_{i-1}})\geq\frac{c_{0}}{T^{1/2}}$
@@
($\mathbb{P}(\{t_{i}-t_{i-1}<T\text{ }\forall L+1\leq i\leq 2L\}\cap\mathcal{G}_{2L-1})\\ =\mathbb{E}\Big{[}\mathbbm{1}_{\{t_{i}-t_{i-1}<T\text{ }\forall L+1\leq i\leq 2L-1\}\cap\mathcal{G}_{2L-1}}\big{(}1-\mathbb{P}(t_{2L}-t_{2L-1}\geq T|\mathcal{F}_{t_{2L-1}})\big{)}\Big{]}.$ (36)
Thanks to Proposition 2.4 we know that the expression on the right-hand side of (36) is at most
@@
Iterating and using the classical inequality $1+x\leq e^{x}$ (which is valid for all $x\in\mathbb{R}$) together with the fact that $\mathcal{G}_{j}\subset\mathcal{G}_{j-1}$ for each $j$, we conclude that